OpenStudy (anonymous):

Solve the system: 7x=y+9 5x^2-y= 8 + 3x

OpenStudy (mathstudent55):

You can use the substitution method. Solve the first equation for y, and substitute what y equals into the second equation.

OpenStudy (anonymous):

this is what i have so far: 7x = y+ 9 -> y = 7x- 9 5x^2 - y = 8 + 3x -> y = 5x^2 - 3x -8 7x- 9 = 5x^2 - 3x -8 =5x^2 - 3x - 8

OpenStudy (anonymous):

form eq 1 find y as y=2x-9 now put this value in 2nd eq. just plug in the value y

OpenStudy (anonymous):

yes its correct step...now move ahead..

OpenStudy (anonymous):

factorize it if possible otherwise use quadratic formula..

OpenStudy (anonymous):

so will it be 5x^2 - 10x +1?

OpenStudy (mathstudent55):

This is correct. 7x- 9 = 5x^2 - 3x -8 Now collect like terms and set it equal to zero. Then solve the quadratic for x.

OpenStudy (mathstudent55):

Yes, it's what you have just above, but set it equal to zero. Solve that quadratic for x. I suggest the quadratic formula.

OpenStudy (anonymous):

i dont really know what to do at this point :$ lol

OpenStudy (mathstudent55):

Have you heard of the quadratic formula?

OpenStudy (anonymous):

yes. so i just plug that in and i have my final answer?

OpenStudy (anonymous):

7x- 9 = 5x^2 - 3x -8 =5x^2 - 3x - 8 am i missing something right here? where the left side go... subtract 7x get 5x^2-10x+1 now you use the quadratic formula for factor it here a=5, b=-10 and c=1 so x=[-b+/-sqrt(b^2-4ac)]/2a x=[10+/-sqrt((-10)^2-4(5)(1)]/2(5) x=[10+/-sqrt(100-20)]/10 x=[10+/-sqrt(80)]/10 x=(10+/-8.944)/10 x=(10+8.944)/10=1.8944 x=(10-8.944)/10=0.1056 x=1.8944 or 0.1056

OpenStudy (mathstudent55):

\(x=\dfrac{10 \pm \sqrt{80}}{10}\) \(x=\dfrac{10 \pm 4\sqrt{5}}{10}\) \(x=\dfrac{5 \pm 2\sqrt{5}}{5}\) \(x=1 + \dfrac{2\sqrt{5}}{5}\) or \(x=1 - \dfrac{2\sqrt{5}}{5}\) So far so good. Now substitute each of those x values into one of the original equations to find the y-values.

OpenStudy (anonymous):

ah thanks guys! i really appreciate all your help:)

OpenStudy (mathstudent55):

You're welcome.